# Effective delocalization in the one-dimensional Anderson model with stealthy disorder

### v1

@@ Abstract
We study analytically and numerically the Anderson model in one dimension with “stealthy” disorder, defined as having a power spectrum that vanishes in a continuous band of wave numbers. Motivated by recent studies on the optical transparency properties of stealthy hyperuniform layered media, we compute the localization length via the perturbation theory expansion of the self-energy. We find that, for fixed energy and small but finite disorder strength, there exists for any finite length system a range of stealthiness $\chi$ for which the localization length exceeds the system size. This kind of “effective delocalization” is the result of the novel kind of correlated disorder that spans a continuous range of length scales, a defining characteristic of stealthy systems. Moreover, we support our analytical results with numerical simulations. Our results may serve as a first step in investigating the role of stealthy disorder in quantum systems, which is of both theoretical and experimental relevance.
@@ I Introduction
Since the seminal work by P. W. Anderson in the late 1950s [1], the interplay between randomness and quantum effects in non-interacting systems has received considerable attention in the condensed matter community [2]. The impact of space dimensionality has been addressed in the scaling theory of localization [3] and in recent investigations [4, 5, 6], including the infinite-dimensional limit [7, 8, 9, 10, 11, 12, 13], relevant for the interacting case, known as many-body localization [14, 15, 16, 17, 18, 19]. In one and two dimensions and for a random uncorrelated onsite potential, it is well known [3] that any small amount of disorder (whose strength is parametrized by its variance $W^{2}$) leads to the localization of the wave functions. In 1D, the localization length $\xi$ at small disorder is proportional to $W^{-2}$ [20].
@@
When introducing correlations in the potential energy, the scenario can change and show interesting new phenomena. This occurs, for instance, in the case of quasi-periodic systems [21, 22], displaying extended, critical, and localized phases, separated by non-trivial mobility edges [23, 24, 25]. Another example, relevant to polymer physics, is that of random dimers [26, 27], where a particle can become delocalized in one dimension if two uncorrelated random energies are assigned at random to pairs of lattice sites. However, only two states are extended in this case. A further interesting case that has received attention in the literature, also because of its experimental relevance in cold-atom experiments [28, 29], is that of speckle potentials, in which the two-point correlation function in momentum space has a triangular form with a high-momentum cutoff. In one-dimensional (1D) systems, speckle correlations suppress back-scattering and can lead to an effective suppression of Anderson localization [30, 31, 32]. The effect of correlated disorder on Anderson localization has also been studied previously in a series of papers [33, 34, 35], where it has been shown that the inverse localization length is proportional to the product of the Fourier transform of the correlation of the disorder and of $W^{2}$ [33]. According to these results, the localization length should diverge if the power spectrum vanishes. However, these results should be regarded as a perturbative outcome including only the first-order term at small $W$. Higher-order terms thus lead to a finite localization length.
@@
A special type of disorder is provided by hyperuniform systems [36, 37], which are characterized by an anomalous suppression of large-scale density fluctuations. In momentum space, they are characterized by a structure factor $\mathcal{S}(k)$ that vanishes as $k\to 0$. Since the first studies in the early 2000s, hyperuniform systems have appeared in a wide variety of systems, ranging from number theory [38, 39] to biological systems [40, 41] (see Ref. [37] for a review).
@@
Among hyperuniform disordered systems, there is a special class that is referred to as stealthy hyperuniformity [42, 43, 44, 45]. In this context, stealthy means that the structure factor of such systems vanishes for a continuous range of wave numbers. Therefore, a stealthy hyperuniform system is such that
@@
It is useful to introduce the stealthiness parameter $0\leq\chi\leq 1/2$, defined as the relative fraction of constrained wave vectors, with $\chi=1/2$ corresponding to the lattice. For the specific case 1D systems, relevant in this paper, $\chi=k_{0}/(2\pi)$ [43]. Stealthy hyperuniform systems display remarkable properties, including superior transport [46], elasticity [47], and wave propagation properties [48, 49] compared to non-stealthy and non-hyperuniform isotropic amorphous states of matter [37].
@@
In a recent study, Klatt et al. [50] studied the electromagnetic wave transmission in disordered stealthy hyperuniform layered media formed by slabs of alternating high and low dielectric constant. By computing the Lyapunov exponent using transfer matrix methods, they found no evidence of localization for system sizes up to 10000 slabs, similar to what is found in periodic systems. From the analytical side, it has been shown in Ref. [48] that the nonlocal strong contrast expansion [51] truncated at third-order predicts that 1D disordered stealthy hyperuniform layered media have perfect transparency (no Anderson localization) for a range of frequencies. It should be mentioned that the nonlocal strong contrast expansion is characterized by fast convergence, and lower-order terms provide accurate approximations for the effective dielectric constant.
@@
Motivated by the aforementioned effects of correlated disorder on Anderson localization and the recent results on the effect of stealthy hyperuniformity on 1D layered media, in this work, we study the effect of stealthy disorder in the 1D Anderson model. In this context, the stealthiness parameter $\chi$ measures the fraction of harmonics for which the two-point correlation function of the random potential vanishes (see Eq. (16)). Unlike the quasi-periodic models, in this case, the system is disordered and allows a continuous range of energies, unlike the random dimer model. It should be mentioned that the effect on localization of non-stealthy hyperuniform disorder in quantum systems has been addressed in the literature, and no significant differences with usual disorder have been reported [52, 53, 54]. However, the study of stealthy hyperuniform disorder has not received attention to the best of our knowledge.
@@
We determine the localization length using a perturbation theory expansion of the self-energy at small disorder $W\ll 1$ for the special case of stealthy hyperuniform disorder. Our procedure can be applied to any kind of stealthy disorder. We are able to determine, for each energy and size of the stealthy region, the power of $W$ responsible for the leading contribution to the localization length. In particular, we find that, for any given length of the 1D system $L$, energy and weak disorder $W\ll 1$, there exists a value of the stealthiness parameter $\chi$ in which the system is disordered but with localization length exceeding the system size, and thus effectively delocalized. If the thermodynamic limit $L\to\infty$ is taken first, the localization length is always finite, thus not violating localization theorems [55, 3].
@@ III Perturbation theory
We are now interested in calculating the dependence of the localization length $\xi(k)$ on $W$ at a given momentum (or, equivalently, energy through the dispersion relation in Eq. (1)). In 1D systems, the localization length is proportional to the mean free path $\ell(k)$, according to $\xi(k)\simeq 2\ell(k)$ [20], which in turn is proportional to the mean survival time $\tau(k)$, defined as $\ell(k)=(\partial_{k}\epsilon_{k})\tau(k)$  [^1] The mean survival time $\tau$ can be interpreted as the average time between two scattering events of the quantum particle on the disordered potential. At this point, it is well known [57] that the mean survival time is related to the imaginary part of the on-shell self-energy $\Sigma(k,\epsilon_{k})$ according to
@@
Following the standard textbook approach [57, 58], we can expand the propagator in powers of $V$ as
@@
For concreteness, we specify a particular form for the stealthy potential, which we choose to be stealthy hyperuniform, motivated by the connection we discussed earlier with optical transparency in layered media [50]. We thus define, restricting to the first Brillouin zone,
@@
where the stealthy interval size $k_{0}$ is related to the $\chi$ parameter commonly used in stealthy hyperuniformity via $\chi=k_{0}/(2\pi)$. Notice that for $\chi=0$, $w_{q}$ are independently distributed Gaussian random variables, and so are the $w_{i}$, recovering the totally uncorrelated case. For $\chi=1/2$, $w_{i}=0$ for all $i$, corresponding to no disorder. The spectrum of the Hamiltonian $H$ with $W=0$ is symmetric and double degenerate; therefore, we can focus on the interval $0\leq k<\pi/2$ without loss of generality.
@@
A few comments are in order. The result in Eq. (21) is well known, and coincides with the finding of Izrailev and Krokhin [33]. This result can also be derived from Boltzmann conductivity [60], and using Fermi’s Golden Rule [61, 62].
@@
The first constraint comes from the stealthy hyperuniform disorder, and the second constraint is given by the perturbation theory order. Clearly, there is a solution only for $q_{1}>0$ and $q_{2}<0$ or vice versa. Such an equation can be satisfied only if $k_{0}<\pi-k$. It is easy to show that $\Sigma^{(2)}_{\mathrm{non-cross}}(k,\epsilon_{k})$ gives the same conditions. Henceforth, one can compute explicitly the correction to $\xi(k)$ at second order, which will be of the form $\xi\sim W^{-4}$.
@@
Figure 3: Dependence of the localization length $\xi(k,k_{0})$ on the disorder strength $W$ at the leading non-vanishing order in perturbation theory. We also indicate the expression for the separatrixes between different regions, whose general expression is given in Eq. (25).
Different regions represent different leading order non-zero contributions to the localization length, indicated by the power in the white circles. The scaling with $W^{-2}$ is the usual Anderson localization behavior that one gets also with ordinary (non-stealthy) disorder. The vertical line at $k_{0}=\chi=0$, corresponding to uncorrelated disorder, always exhibits this behavior, as expected.
@@
The series in Eq. (26) is an asymptotic expansion in small $W^{2}$, which is diverging when non-perturbative effects are excluded, as very often occurs in perturbative calculations [63]. Once such non-perturbative effects are taken into account, the divergence is removed and the result can be written as a finite number of terms up to $n=N$ plus a small remainder, which is of order $o(W^{2N})$. When doing so, the vanishing of the lower order terms due to the stealthiness condition implies a larger localization length, scaling with the first non-zero term in $W^{2}$. In the next Section, we will show that this is indeed the case by numerically computing the dependence of the localization length $\xi$ on $W$ at different values of $\chi$, which properly accounts for the remainder terms.
@@ IV Numerical results
which corresponds to the usual definition of fractal dimension $D_{1}$ found in the literature [64, 65, 66].
@@
We show in Fig. 4 the localization length as a function of $\chi$ for a few values of $W$, where the energy is fixed at E=1.8E=-1.8, corresponding to the red dashed line in Fig. 3. By increasing $\chi$, the localization length $\xi$ shows sudden increases. Notice that, by decreasing $W$, the transitions in numerical curves are becoming steeper and the $\chi$ values where they occur are going towards the perturbation theory prediction, reported as vertical black dashed lines, corresponding to the first two red dots from the left in Fig. 3. Subsequent transitions are not visible due to the limited system sizes that can be reached numerically.
@@ V Discussion
We can summarize our results as follows. For an arbitrarily large system size $L$, fixed small disorder $W$, and fixed energy $E$, there exists a stealthy interval size $\chi<1/2$ for which $a_{2n}=0$ (see Eq. (26)) for all $n<m$ with $m$ arbitrarily large, thus meaning that the localization length $\xi$ exceeds the system size $L$. Therefore, for all practical purposes, there is a finite continuous range of effectively extended states, associated with the energy $E$, and the system is effectively delocalized.
@@
We obtained the aforementioned result by computing the localization length in the 1D Anderson model with stealthy disorder, such that its power spectrum vanishes for a continuous band. From the perturbative expansion of the imaginary part of the self-energy, we were able to determine, for each value of the energy and of the stealthiness parameter $\chi$, the leading contribution $W^{-2n}$, $n\geq 1$, to the localization length $\xi$, summarized in Fig. 3. We supported our findings with a numerical analysis of the localization length extracted from the fractal dimension of eigenstates of a given energy. Our results show that stealthy disorder provides a way to adjust the scaling of the localization length and to make it bigger than the system size without taking the limit of zero disorder ($W\to 0$).
Disordered stealthy hyperuniform materials exhibit a unique blend of liquid-like and crystal-like characteristics: they lack Bragg peaks yet prevent the formation of arbitrarily large voids in their microstructure. Our results suggest that stealthy disorder can be viewed as an intermediate regime also in the context of localization in the realm of quantum disordered systems, bridging the phenomena observed in ordinary uncorrelated random potentials and more finely tuned correlated structures, such as quasi-periodic potentials or even the absence of potential energy altogether. The stealthiness parameter $\chi$ serves as a control knob, allowing for continuous setting of the degree of stealthiness across this spectrum.
Our findings lend support to the results of Klatt et al. [50], although they pertain to two different models. The absence of single back-scattering for $k<\pi\chi$ and contributions to the localization length proportional to $W^{-2n}$, $n\geq 2$, is analogous to the complete transparency found in Ref. [50] up to third order in the strong contrast expansion.
@@
From a practical perspective, the Anderson model with stealthy disorder could be realized on currently available experimental platforms, such as ultracold atoms in optical lattices or programmable quantum simulators, which provide a high degree of tunability of the onsite potential [29, 67, 68, 69].
The role of stealthy disorder in interacting quantum systems and in higher dimensions is an intriguing question that we leave for future work.

### v3

@@ Abstract
We study analytically and numerically the Anderson model in one dimension with “stealthy” disorder, defined as having a power spectrum that vanishes in a continuous band of wave numbers. Motivated by recent studies on the optical transparency properties of stealthy hyperuniform layered media, we compute the localization length using a perturbative expansion of the self-energy. We find that, for fixed energy and small but finite disorder strength $W$, there exists for any finite length system a range of stealthiness $\chi$ for which the localization length exceeds the system size. This kind of “effective delocalization” is the result of the novel kind of correlated disorder that spans a continuous range of length scales, a defining characteristic of stealthy systems. Unlike uncorrelated disorder, for which the localization length $\xi$ scales as $W^{-2}$ to leading order for small W, the leading order terms in the perturbation expansion of $\xi$ for stealthy disordered systems vanish identically for a progressively large number of terms as $\chi$ increases such that $\xi$ scales as $W^{-2n}$ with arbitrarily large $n$. Moreover, we support our analytical results with numerical simulations. Our results introduce stealthy disorder into quantum tight-binding models and show that enforcing a low-$k$ spectral gap markedly alters the scattering landscape, enabling localization lengths that exceed the system size at fixed disorder strength. Since this mechanism relies only on the spectral properties of the disorder, it carries over directly to photonic and phononic wave systems.
@@ I Introduction
Since the seminal work by P. W. Anderson in the late 1950s [Anderson1958absence], the interplay between randomness and quantum effects in non-interacting systems has received considerable attention in the condensed matter community [evers2008anderson]. The impact of space dimensionality has been addressed in the scaling theory of localization [abrahams1979scaling] and in recent investigations [ueoka2014dimensional, Tarquini2017critical, altshuler2025Renormalization], including the infinite-dimensional limit [tikhonov2016Anderson, bera2018return, sierant2023universality, vanoni2023renormalization, Derrida1980REM, Baldwin2016ManyBody, Balducci2025Scaling], relevant for the interacting case, known as many-body localization [altshuler1997quasiparticle, Basko06, oganesyan2007localization, ros2015integrals, tikhonov2021AndersonMBL, niedda2024renormalizationgroupanalysismanybodylocalization]. In one and two dimensions and for a random uncorrelated onsite potential, it is well known [abrahams1979scaling] that any small amount of disorder (whose strength $W^{2}$ is parametrized by its variance) leads to the localization of the wave functions. In 1D, the localization length $\xi$ at small disorder is proportional to $W^{-2}$ [Lee1985Disordered].
@@
When introducing correlations in the potential energy, the scenario can change and show interesting new phenomena. This occurs, for instance, in the case of quasi-periodic systems [Aubry1980Annals, Luschen2018], displaying extended, critical, and localized phases, separated by non-trivial mobility edges [Ganeshan2015Nearest, Biddle2010Predicted, Goncalves2023Critical]. Another example, relevant to polymer physics, is that of random dimers [Dunlap1990Absence, Phillips1991Localization], where a particle can become delocalized in one dimension if two uncorrelated random energies are assigned at random to pairs of lattice sites. However, only two states are extended in this case. A further interesting case that has received attention in the literature, also because of its experimental relevance in cold-atom  [Sanchez2007Anderson, Billy2008] and photonics [schwartz2007transport, dikopoltsev2022observation] experiments, is that of speckle potentials, in which the two-point correlation function has a high-momentum cutoff. In one-dimensional (1D) systems, speckle correlations suppress back-scattering and can lead to an effective suppression of Anderson localization [Lugan2009Onedimensional, Piraud2012Anderson, morpurgo2025localizationtransitioninteractingquantum]. The effect of correlated disorder on Anderson localization has also been studied previously in a series of papers [izrailev_localization_1999, izrailev_anomalous_2012, Izrailev1995Hamiltonian], where it has been shown that the inverse localization length is proportional to the product of the Fourier transform of the correlation of the disorder and $W^{2}$ [izrailev_localization_1999]. According to these results, the localization length should diverge if the power spectrum vanishes. However, these results should be regarded as a perturbative outcome including only the first-order term at small $W$. Higher-order terms thus lead to a finite localization length.
@@
A special type of disorder is provided by hyperuniform systems [To03a, To18a], which are characterized by an anomalous suppression of large-scale density fluctuations. In momentum space, they are characterized by a structure factor $\mathcal{S}(k)$ that vanishes as $k\to 0$. Since the first studies in the early 2000s, hyperuniform systems have appeared in a wide variety of systems, ranging from number theory [To18d, To08b] to biological systems [jiao_avian_2014, Ma15] (see Ref. [To18a] for a review).
@@
Among hyperuniform disordered systems, there is a special class that is referred to as stealthy hyperuniformity [uche_constraints_2004, batten_classical_2008, Zh15a, Zh15b]. In this context, stealthy means that the structure factor of such systems vanishes for a continuous range of wave numbers. Therefore, a stealthy hyperuniform system is such that
@@
It is useful to introduce the stealthiness parameter $0\leq\chi\leq 1/2$, defined as the relative fraction of constrained wavevectors. For point patterns, the configuration converges to a periodic lattice in the limit $\chi\to 1/2$. For the specific case 1D systems, relevant in this paper, $\chi=k_{0}/(2\pi)$ [batten_classical_2008]. Stealthy hyperuniform systems display remarkable properties, including superior transport [To21d], elasticity [kim_effective_2020], and wave propagation properties [kim_effective_2023, Vanoni2025Dynamical] compared to non-stealthy and non-hyperuniform isotropic amorphous states of matter [To18a].
@@
In a recent study, Klatt et al. [klatt2025transparencyversusandersonlocalization] studied the electromagnetic wave transmission in disordered stealthy hyperuniform layered media formed by slabs of alternating high and low dielectric constant. By computing the Lyapunov exponent using transfer matrix methods, they found no evidence of localization for system sizes up to 10000 slabs, similar to what is found in periodic systems. On the analytical side, it has been shown in Ref. [kim_effective_2023] that the nonlocal strong contrast expansion [torquato_nonlocal_2021] truncated at third-order predicts that 1D disordered stealthy hyperuniform layered media have perfect transparency (no Anderson localization) for a range of frequencies. Therefore, Anderson localization could only arise at fourth order and beyond. It should be mentioned that the nonlocal strong contrast expansion is characterized by fast convergence, and lower-order terms provide accurate approximations for the effective dielectric constant.
@@
Motivated by the aforementioned effects of correlated disorder on Anderson localization and the recent results on the effect of stealthy hyperuniformity on 1D layered media, in this work, we study the effect of stealthy disorder in the 1D Anderson model. In this context, the stealthiness parameter $\chi$ measures the fraction of harmonics for which the power spectrum of the random potential vanishes (see Eq. (16)). Unlike the quasi-periodic models, in this case, the system is disordered and allows a continuous range of energies, unlike the random dimer model. It should be mentioned that the effect on localization of non-stealthy hyperuniform disorder in quantum systems has been addressed in the literature [Crowley2019Quantum, Shi2022Manybody]; a recent finding, for example, showed that Lifshitz tails have different scalings compared to uncorrelated disorder, effectively widening the band gap [Karcher2024Effect]. Experiments on disordered two-dimensional stealthy photonic crystal systems have demonstrated non-trivial wave transport properties with both strong disorder and non-Hermiticity [barsukova2025stealthy]. However, the latter work did not examine localization behavior.
@@
We determine the localization length using a perturbation theory expansion of the self-energy at small disorder $W\ll 1$ for the special case of stealthy hyperuniform disorder. Our procedure can be applied to any kind of stealthy disorder. We are able to determine, for each energy and size of the stealthy region, the power of $W$ responsible for the leading contribution to the localization length. In particular, we find that, for any given length of the 1D system $L$, energy and weak disorder $W\ll 1$, there exists a value of the stealthiness parameter $\chi$ in which the system is disordered but with localization length exceeding the system size, and thus effectively delocalized. If the thermodynamic limit $L\to\infty$ is taken first, the localization length is always finite, thus not violating localization theorems [Wegner1976, abrahams1979scaling].
@@ III Perturbation theory
We are now interested in calculating the dependence of the localization length $\xi(k)$ on $W$ at a given momentum (or, equivalently, energy through the dispersion relation in Eq. (1)). In 1D systems, the localization length is proportional to the mean free path $\ell(k)$, according to $\xi(k)\simeq 2\ell(k)$ [Lee1985Disordered], which in turn is proportional to the mean survival time $\tau(k)$, defined as $\ell(k)=(\partial_{k}\epsilon_{k})\tau(k)$  [^1] The mean survival time $\tau$ can be interpreted as the average time between two scattering events of the quantum particle on the disordered potential. At this point, it is well known [Fetter2003Quantum] that the mean survival time is related to the imaginary part of the on-shell self-energy $\Sigma(k,\epsilon_{k})$ according to
@@
Following the standard textbook approach [Fetter2003Quantum, Akkermans_Montambaux_2007], we can expand the propagator in powers of $V$ as
@@
For concreteness, we specify a particular form for the stealthy potential, which we choose to be stealthy hyperuniform, motivated by the connection we discussed earlier with optical transparency in layered media [klatt2025transparencyversusandersonlocalization]. We thus define, restricting to the first Brillouin zone,
@@
where the stealthy interval size $k_{0}$ is related to the $\chi$ parameter commonly used in stealthy hyperuniformity via $\chi=k_{0}/(2\pi)$. Notice that for $\chi=0$, $w_{q}$ are independently distributed Gaussian random variables, and so are the $w_{i}$, recovering the totally uncorrelated case. For $\chi=1/2$, $w_{i}=(-1)^{i}$, corresponding to no disorder. The spectrum of the Hamiltonian $H$ with $W=0$ is symmetric and double degenerate; therefore, we can focus on the interval $0\leq k<\pi/2$ without loss of generality.
@@
A few comments are in order. The result in Eq. (21) is well known, and coincides with the finding of Izrailev and Krokhin [izrailev_localization_1999]. This result can also be derived from Boltzmann conductivity [herbut2000commentlocalizationmobilityedge], and using Fermi’s Golden Rule [Licciardello_1975_Conductivity, scardicchio2017perturbationtheoryapproachesanderson]. Note also that the expression in Eq. (21) is analogous to the one found within the strong-contrast expansion for the dynamic dielectric constant in 1D stealthy hyperuniform two-phase media [kim_effective_2023], where the power spectrum at momentum $2k$ is replaced by the spectral density at momentum $2k$.
@@
The first constraint comes from the stealthy hyperuniform disorder, and the second constraint is given by the perturbation theory order. Clearly, there is a solution only for $q_{1}>0$ and $q_{2}<0$ or vice versa. Such an equation can be satisfied only if $k_{0}<\pi-k$. It is easy to show that $\Sigma^{(2)}_{\mathrm{non-cross}}(k,\epsilon_{k})$ gives the same conditions. Henceforth, one can compute explicitly the correction to $\xi(k)$ at second order, which will be of the form $\xi\sim W^{-4}$. We report the calculation in the Supplemental Material.
@@
Figure 3: Phase diagram showing the dependence of the localization length $\xi(k,k_{0})$ on the disorder strength $W$ at the leading non-vanishing order in perturbation theory. We also indicate the expression for the separatrices between different regions, whose general expression is given in Eq. (25). The red dotted line indicates the energy for which we will present numerical results.
Different regions represent different leading order non-zero contributions to the localization length, indicated by the power in the white circles. The scaling with $W^{-2}$ is the usual Anderson localization behavior that one also gets with ordinary (non-stealthy) disorder. The vertical line at $k_{0}=\chi=0$, corresponding to uncorrelated disorder, always exhibits this behavior, as expected.
@@
The series in Eq. (26) is an asymptotic expansion in small $W^{2}$, which is diverging when non-perturbative effects are excluded, as very often occurs in perturbative calculations [Dyson1952Divergence]. Once such non-perturbative effects are taken into account, the divergence is removed and the result can be written as a finite number of terms up to $n=N$ plus a small remainder, which is of order $o(W^{2N})$. When doing so, the vanishing of the lower order terms due to the stealthiness condition implies a larger localization length, scaling with the first non-zero term in $W^{2}$. In the next Section, we will show that this is indeed the case by numerically computing the dependence of the localization length $\xi$ on $W$ at different values of $\chi$, which properly accounts for the remainder terms.
In this section, we have discussed the series of transitions in the leading order contribution to the localization length at small disorder $W$. In the Supplementary Material, we show that analogous transitions occur in the level splitting of adjacent energy levels that become degenerate as $W\to 0$ due to the $k\to-k$ symmetry.
@@ IV Numerical results
which corresponds to the usual definition of fractal dimension $D_{1}$ found in the literature [Chalker1996Spectral, Bogomolny2011Eigenfunction, kutlin2024investigating].
@@
We show in Fig. 4 the localization length as a function of $\chi$ for a few values of $W$, where the energy is fixed at E=1.8E=-1.8, corresponding to the red dashed line in the top panel of Fig. 3. By increasing $\chi$, the localization length $\xi$ shows sudden increases. Notice that, by decreasing $W$, the transitions in numerical curves are becoming steeper and the $\chi$ values where they occur are going towards the perturbation theory prediction, reported as vertical black dashed lines, corresponding to the first two red dots from the left in the top panel of Fig. 3. Subsequent transitions are not visible due to the limited system sizes that can be reached numerically.
@@ V Discussion
Generically, disordered stealthy hyperuniform materials lack Bragg peaks similar to liquid-like disordered systems and yet suppress long-range density fluctuations as crystals. Our results indicate that stealthy disorder represents an intermediate regime when it comes to localization in 1D quantum disordered systems as well. The stealthiness parameter $\chi$ acts as a control knob that varies between $\chi=0$, corresponding to a random potential whose values are uncorrelated, and $\chi=1/2$, corresponding to a periodic potential. Stealthy disorder leads to effective delocalization, in which the localization length scales as $W^{-2n}$ where $n>1$.
@@
More precisely, for an arbitrarily large system size $L$, fixed small disorder $W$, and fixed energy $E$, there exists a stealthy interval size $\chi<1/2$ for which $a_{2n}=0$ (see Eq. (26)) for all $n<m$ with $m$ arbitrarily large, thus meaning that the localization length $\xi$ exceeds the system size $L$. Therefore, for all practical purposes, there is a finite continuous range of effectively extended states, associated with the energy $E$, and the system is effectively delocalized. In the limit $\chi\rightarrow 1/2$, we expect non-perturbative effects not captured by our treatment to make non-negligible contributions. We leave the investigation of this region of the phase diagram for future work.
@@
We obtained the aforementioned result by computing the localization length in the 1D Anderson model with stealthy disorder, such that its power spectrum vanishes for a continuous band. From the perturbative expansion of the imaginary part of the self-energy, we were able to determine, for each value of the energy and of the stealthiness parameter $\chi$, the leading contribution $W^{-2n}$, $n\geq 1$, to the localization length $\xi$, summarized in Fig. 3. We supported our findings with a numerical analysis of the localization length extracted from the fractal dimension of eigenstates of a given energy. Our results show that stealthy disorder makes it possible to tune how the localization length scales, enabling it to become larger than the system size even when the disorder variance is held fixed. Unlike uncorrelated disorder, for which $\xi$ scales as $W^{-2}$ to leading order for small $W$, the result for stealthy disordered systems is that the leading order terms in the perturbation expansion of $\xi$ vanish identically for a progressively large number of terms as $\chi$ increases such that $\xi$ scales as $W^{-2n}$ with arbitrarily large $n$, where $n$ is a positive integer.
Our findings for the Anderson model are reminiscent of recent studies on the propagation of light in a one-dimensional layered medium [kim_effective_2023, torquato_nonlocal_2021, klatt2025transparencyversusandersonlocalization]. The disorder strength $W$ in the Anderson problem is analogous to the dielectric contrast in the two-phase layered medium [torquato_nonlocal_2021]. In the latter case, complete transparency can be shown to occur at least up to third order in the strong dielectric contrast expansion [kim_effective_2023, torquato_nonlocal_2021] for a finite continuous range of frequencies, which matches the result for the tight-binding Anderson model where contributions to the localization length are zero to $2n$-th order for arbitrarily large finite $n$. Moreover, the momentum dependence of the imaginary part of the self-energy in the Anderson problem parallels that of the nonlocal attenuation function in the 1D stealthy two-phase medium [kim_effective_2023].
@@
From a practical perspective, the Anderson model with stealthy disorder can be realized on currently available experimental platforms, such as ultracold atoms in optical lattices, waveguide arrays, photonic lattices in photorefractive crystals in the paraxial regime of propagation, or other programmable quantum simulators, which provide a high degree of tunability of the potential [Billy2008, roati2008anderson, Sc07b, lahini_anderson_2008, Schreiber2015Observation, Smith2016]. In addition, stealthy hyperuniformity was recently realized and probed experimentally in a two-dimensional photonic crystal slab platform [barsukova2025stealthy]. We leave the intriguing question of stealthy hyperuniformity in interacting quantum systems for future work.